Shift report 20010905 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Wed Sep 05 2001 - 16:03:20 EDT

  • Next message: Apache: "Shift report 20010906 16:00-24:00"

    ------------------------------------------------------------
    Date: Wed Sep  5 16:03:20 2001, Shift: 08:00-16:00
    Supervisor: K. Hagel, with: 
    Log book#8 pages 79 - 80
    
    Shift Summary:
    ------------------------------------------------------------
    No beam during entire shift.  Toward end of shift, 
    machine was at injection; then a power supply problem.
    
    Spent a lot of time with DB/FV working on ZDCs.
    1. Changed one BNC-LEMO converter on left 1 which 
       was injecting noise into both it's signal as well
       as the sum signal
    2. Changed two filters as they were of an obviously 
       different vintage than the ones fabricated for the
       ZDCs.  This eliminated an overshoot problem (at 
       least with the ZDC LED) and made the analog sum
       agree much closer with the digital sum.  Currently
       analog Sum/digital Sum Left  = .85
       analog Sum/digital Sum Right = .94
       For the left, the ratio of .85 is up from .4 before we
       started the diagnosis
    3. Tried to change the pedestal in right 1.  I tried many
       iterations with the ADC on the CAMAC extender, but I
       could never make a pedestal appear.  Conclude that 
       pedestal adjustment of chan 4 of the ADC does not work.
       The LED signal comes in OK, however.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Wed Sep 05 2001 - 16:03:56 EDT